OpenStudy (anonymous):

i don't know the steps to solving this equation :3 Solve the following inequalities, and graph your solution on a number line: a. 6x^2+29x >-30 b. x^2+ 3x + 4 > 0

OpenStudy (anonymous):

Hint: Use quadratic formula

OpenStudy (anonymous):

so you just plug in the numbers and you're done?

OpenStudy (anonymous):

a.6x^2+29x >-30 <=>6x^2+29x+30 >0 <=>x=-3/2 or x=-10/3 if x <-3/2, the other x must be < -10/3 for the ans to be 0 if x > -3/2, the other x must be > -10/3 for the ans to be 0 =>x>-3/2 => x<-10/3 just do the same thing with the second one, then you'll have the solutions

OpenStudy (anonymous):

im still confused...

OpenStudy (anonymous):

which part?